You are taking some time off from studying and decide to go to the beach. It is a hot August day so you decide to bring a Styrofoam cooler. This particular cooler is made with outside dimensions of 3.00 × 2.00 × 1.50 ft. The Styrofoam is 0.25 ft thick on all sides including the lid and has a density of 50 kg/m3.a.
a. If you can comfortably lift 30 lb, is this a good cooler to bring to the beach?
b. If the cooler is filled to the top with water at 2ºC, what is the mass of the cooler and water (in kg)?
1
Expert's answer
2019-12-06T09:43:20-0500

a)

"V=2(0.25)((3)(2)+(3)(1.5)+(1.5)(2))=6.75\ ft^3=0.17\ m^3"

"m=(50)(0.17)=8.5\ kg=3.6\ lb"

It il less than 30 lb.


b)

"V'=(2)(3)(1.5)-6.75=2.25\ lb^3=0.0566\ m^3"


"M=8.5+(0.0566)1000=65.1\ kg"